### v1

# Preference Distillation for Personalized Generative Recommendation

Abstract
Recently, researchers have investigated the capabilities of Large Language Models (LLMs) for generative recommender systems. Existing LLM-based recommender models are trained by adding user and item IDs to a discrete prompt template. However, the disconnect between IDs and natural language makes it difficult for the LLM to learn the relationship between users. To address this issue, we propose a PErsonAlized PrOmpt Distillation (PeaPOD) approach, to distill user preferences as personalized soft prompts. Considering the complexities of user preferences in the real world, we maintain a shared set of learnable prompts that are dynamically weighted based on the user’s interests to construct the user-personalized prompt in a compositional manner. Experimental results on three real-world datasets demonstrate the effectiveness of our PeaPOD model on sequential recommendation, top-n recommendation, and explanation generation tasks.
Preference Distillation for Personalized Generative Recommendation
Jerome Ramos  Bin Wu  Aldo Lipani University College London, London, UK {jerome.ramos.20, bin.wu.23, aldo.lipani}@ucl.ac.uk
1 Introduction
Recommender systems Ge et al. (2022) have emerged as pivotal tools in guiding users through vast amounts of items and tailoring personalized experiences across various tasks, including sequential recommendation, top-n recommendation and explanation generation. Recently, recommender systems trained using large language models (LLMs), known as generative recommender systems, have unlocked the ability to perform multiple tasks with a single model. It has been shown to perform well in a zero- and few-shot setting Sanner et al. (2023) and can be further fine-tuned on a specific domain to improve performance  Ramos et al. (2024).
P5  Geng et al. (2022) formulates recommendation as a text-to-text sequence task and trains the model to perform generative recommendation based on discrete prompt templates. User and item IDs serve as important identifiers for discrete prompt templates, beyond the simple semantic meaning within their word embeddings. Thus, the disconnect between IDs and natural language and the inherent rigidity of discrete prompts makes it difficult for the LLM to learn the relationship between users, limiting the effectiveness of this approach. Prompt Distillation has been proposed to distill the knowledge in discrete prompts into continuous prompt vectors which do not map to any specific words Li et al. (2023). However, the existing work distills the knowledge into global-shared prompts, which is in conflict with the fact that users’ personalized preferences greatly differ across the entire user population. Given that each user has their own unique preferences (Wu et al., 2022), it is imperative to design a method to effectively distill preferences for personalized recommendation.
One intuitive way to introduce distilled preferences is to create unique prompts per user based solely on their individual interaction history. However, there are two key issues: (1) Distilling separated prompts for each user is costly and difficult, where the number of prompts increases with the number of the user and in most cases, each user’s interaction history is too sparse to effectively personalize the recommender; and (2) This approach does not allow the model to match users with similar interests, which has been shown to increase performance for traditional recommender models in a warm-start setting Koren et al. (2009); Mnih and Salakhutdinov (2007); Cheng et al. (2016).
In this work, we propose PErsonAlized PrOmpt Distillation (PeaPOD), a more flexible and effective framework to distill user preferences as personalized soft prompts. Specifically, we construct a set of decomposed prompt components, which maintains a many-to-many mapping between user preferences and continuous prompts. This is due to the compositional nature of user preferences, meaning that users might share similar interests but maintain their distinction as well. By updating only a small set of prompt components, we can efficiently share prompts among users rather than maintaining a unique prompt per user.
Moreover, to construct the many-to-many mapping (Wu et al., 2023) between users and the distilled prompt components, we propose a compositional strategy to dynamically weight the prompts based on user preference. First, we use matrix factorization Koren et al. (2009) to generate a user embedding, which compresses a user’s interaction history into a high-dimensional space. The user embeddings are used to identify user preferences, helping to create a weighted summation of the decomposed prompt components, called the user-personalized prompt. Importantly, the construction of the user-personalized prompt are dynamic across users but only maintains a limited set of shared prompt components. From experiments across three datasets, PeaPOD is able to achieve state-of-the-art performance in three recommendation tasks. We provide all data and code used at: https://github.com/jeromeramos70/peapod.
In summary, our key contributions are:
  1. 1.
    We provide a novel architecture, namely PErsonAlized PrOmpt Distillation (PeaPOD), to distill complex user preferences into a limited set of learnable prompts.
  2. 2.
    We propose a compositional strategy, dynamically weighting the decomposed prompt components to construct the user-specific prompt based on user preference.
  3. 3.
    We conduct extensive experiments, showing the effectiveness and superiority of our proposed PeaPOD model.
2 Related Works
2.1 LLMs for Recommendation
In recent years, researchers have explored the effectiveness of using LLMs for recommendation. The two main methods of LLMs for recommender systems are (1) using the LLM’s pre-trained knowledge to recommend items in a zero-shot or few-shot setting and (2) fine-tuning the model for a given domain.
One advantage of LLMs is that they are pre-trained with a large amount of data from various domains. Thus,  Sanner et al. (2023) crowdsourced a dataset of natural language summaries of user preferences and showed that prompting a language model with these user summaries yielded competitive performance with traditional baselines in a near cold-start setting.  Ramos et al. (2024) showed that a generative recommendation system trained solely on natural language user profiles are also competitive with traditional recommender systems in a warm-start setting, with the added benefit of transparency and scrutability. In addition, the zero-shot prompting of GPT-4 has also beaten SOTA models in a conversational recommender setting He et al. (2023). Although prompting LLMs can lead to strong performance in recommender tasks, they are heavily reliant on better content knowledge rather than collaborative knowledge He et al. (2023); Wu et al. (2024). Thus, if the item metadata is not already learned in the pre-trained knowledge base or if the target dataset users interact differently than the users that the LLM was originally trained on, the performance of the recommender system will suffer.
To mitigate these issues, the model can be further fine-tuned on target datasets in order to train the model based on the given user population and add or update its knowledge on the item dataset. P5 Geng et al. (2022) is the first work that trains a T5-based Raffel et al. (2020) foundational recommender model to unify various recommendation tasks in a shared framework. VIP5 Geng et al. (2023) extended P5 to a multi-modal foundational recommender system. POD Li et al. (2023) builds upon the P5 model and appends a task-specific, continuous prompt to the input to improve performance. Although these models pass in the user_id as inputs, they still do not leverage the collaborative information contained from similar users. In particular, they do not implement a user-specific component to identify and share knowledge between users with similar preferences. Previous work in continual learning for image classification has shown that an attention-based mechanism based on the image input is able to learn similarities across similar images to achieve state-of-the-art performance Smith et al. (2023). In this work, we focus on adding an attention-based mechanism to our model in order to leverage the shared collaborative knowledge between users to improve the recommendation performance.
2.2 Structural Prompt
One issue with generative approaches to recommendation is how to effectively encode user-item interactions to use with LLMs. Researchers have proposed numerous methods to bridge the gap between item IDs and natural language. One option is to add the user’s name and item title to the input prompt. However, this data may be unavailable or non-unique, making it difficult to personalize recommendations. Another option is to directly insert the user and item IDs to the discrete prompt template. However, LLMs are trained on natural language text rather than IDs, which limits the effectiveness of this approach.
To tackle these issues, many researchers have explored how to effectively encode item IDs with LLMs. For example, Rajput et al. (2023) proposed using RQ-VAE with item metadata to generate semantically meaningful tuples of tokens to serve as the Semantic ID for each item. GPTRec uses a novel SVD tokenization method to generate quantized item embeddings based on user-item interactions Petrov and Macdonald (2023).  Tan et al. (2024) proposed training a textual ID generator alongside the LLM-based recommender to convert item ids to meaningful text descriptions. Although these approaches have improved item IDs encoding for generative recommendations, the effective encoding of user information remains largely unexplored. In particular, existing works do not tackle the problem of how to encode both a user’s personalized preferences and the shared knowledge between similar users to an LLM. Thus, our work focuses on how to develop user-personalized prompts to improve the performance of a generative recommender system.
3 Preliminary
3.1 Recommendation Tasks
Following the existing works Geng et al. (2022), we focus mainly on three important recommendation tasks: (1) Sequential Recommendation, (2) Top-n Recommendation, and (3) Explanation Generation. Let $u\in U$ denote a user in the user set and $i\in I$ denote an item in the item set across all three tasks. We aim to train a unified model to complete the following tasks:
  • Sequential Recommendation: given a user $u$ and a list of historical interactions in chronological order $i_{1},i_{2},...i_{n-1}$, the task is to predict the next item $i_{n}$ that $u$ will interact with.
  • Top-n Recommendation: given a pool of randomly sampled items from $I$, the model must select the top $N$ items that the user wants to interact with.
  • Explanation Generation: given a user $u$ and item $i$, the model must generate an explanation $E_{u,i}$ that explains why $u$ likes/dislikes $i$.
Importantly, the scope of our study focuses only on user and item IDs for recommendation and excludes additional information such as item metadata or user profiles because this information may be sensitive or unavailable.
3.2 Prompt Distillation
In traditional recommender systems, user and item IDs are commonly used to identify unique users and items. However, since these IDs are unique to the dataset and are not natural language text, it is necessary to perform additional fine-tuning to allow the LLM to effectively understand these IDs. The most common methodology is to train the model with discrete prompt templates. For example, in the case of explanation generation, the discrete prompt could be “Explain why user_1234 enjoys item_5678”. When using IDs with discrete prompt templates, recent research proposes adding an additional embedding space, called whole-word embeddings, so that the model can recognize which tokens belong to the appropriate ID Geng et al. (2022). In particular, each ID consists of the same whole-word embedding and all other tokens are tokenized as normal.
Although discrete prompts can lead to good performance, they are inflexible and require manually crafted prompts. Thus, the performance of the model may depend strongly on how well the instructions were written. On the other hand, continuous prompts are dynamically adjusted and fine-tuned through techniques like soft prompt tuning, which optimizes prompts by learning continuous embeddings. This approach allows the model to learn regardless of the the quality of the prompt template. Li et al. (2023) introduced Prompt Distillation, which distills discrete prompts into continuous prompt vectors. The authors used prompt distillation to add task-specific continuous prompts per recommendation task globally shared across all users to improve the model’s performance. Since this method adds a shared continuous prompt per recommendation task, we will refer to these continuous prompts as task-specific prompts. Although task-specific prompts were shown to improve performance in our recommendation tasks, they cannot express personalization because each user uses the exact same prompts. Thus, it is imperative to construct a new method to distill each user’s preferences into a personalized soft prompt.
4 User-Personalized Prompt
Figure 1: Method to create the user-personalized prompt. The user embedding from matrix factorization is used to weight the decomposed prompt components. The distillled user preferences are then compositionally combined into a single soft prompt that can be passed to the model as input.
Figure 2: Full PeaPOD model architecture. PeaPOD prepends two continuous prompts to the discrete prompt template: (1) a user-personalized prompt based on the user’s preferences and (2) a global task-specific prompt shared across all users. We add the additional whole-word embedding to distinguish tokens belonging to a given ID and train an encoder-decoder model.
The goal of the user-personalized prompt is to distill user preferences into a soft prompt that is able to represent the user’s personal interests while also leveraging collaborative knowledge between similar users. We maintain a set of decomposed prompt components, which represents a many-to-many mapping between user preferences and continuous prompt vectors. This allows us to efficiently maintain a limited group of prompts and share knowledge across similar users. The decomposed prompt vectors are then dynamically weighted and combined based on the user’s interaction history to form a single continuous prompt. Using this methodology, user with similar preferences should have similar weight distributions. For example, if a decomposed prompt component contained information about action movies, then all users with an affinity to actions movies will contribute significant weight to this prompt component.
4.1 Query Function
Since we want to weight the decomposed prompt components based on each user’s preferences, we introduce a query function $q(x)$. The purpose of $q(x)$ is to obtain an embedding that effectively captures a user’s personalized interests. In our case, we choose to use a user embedding generated from matrix factorization Koren et al. (2009) as the query function. In particular, given a user-item feedback matrix $A\in\mathbb{R}^{|U|\times|I|}$, the model learns a user embedding $U\in\mathbb{R}^{|U|\times D}$ and item embedding $U\in\mathbb{R}^{|I|\times D}$, where $D$ is the dimension size and $|U|$ and $|I|$ are the number of users and items, respectively. By using a user embedding as the query function, we can effectively encode user preferences into the shared set of prompt components.
4.2 Prompt-Component Weighting
We use an attention-based architecture Vaswani et al. (2017) to attend the prompt components based on the user’s personalized preferences. Let $A\in\mathbb{R}^{M\times D}$ be a learnable attention vector, where $M$ is the number of decomposed prompt components. We calculate the element-wise multiplication between q(x) and $A$ to obtain the attended query, denoted as $\lambda$. The attended query allows the model to focus on the most important aspects of the user’s interaction history. To obtain the attended query we calculate:
$\lambda=q(x)\odot A$ (1)
where $\odot$ is the element-wise product. We then calculate the cosine similarity, denoted as $\gamma$, between $\lambda$ and $K\in\mathbb{R}^{M\times D}$, where $k_{M}\in K$ is a learnable key that has a corresponding attention vector $A_{M}$ and prompt component $p_{M}$. The cosine similarity between the attended query $\lambda$ and $K$ is used to determine the magnitude that each prompt component should receive, which we define as:
$\alpha=\gamma(\lambda,K)$ (2)
where $\alpha\in\mathbb{R}^{M\times D}$ is a weighting vector. Since $\alpha$ is updated during training, the model is able to update its representation of the user embedding rather than fixing it to a single representation. $\alpha$ can then be used to weight the prompt components based on attended query function. Thus, we can determine how much each preference should contribute to the user’s overall user-personalized prompt.
4.3 Prompt Composition
Finally, we introduce the user-personalized prompt $P$, which is calculated as a weighted summation over the set of decomposed prompt components:
$P=\sum_{M}\alpha_{M}p_{M}$ (3)
where $p\in\mathbb{R}^{L\times D\times M}$ is a decomposed prompt component and $L$ is the length of each prompt component. Each decomposed prompt component is a soft prompt that represents information that can be collaboratively shared across similar users. We base this idea on the assumption that users may share similar interests and this shared collaborative knowledge can be used to improve personalization. Our many-to-many mapping between user preferences and soft prompts are dynamically weighted using $\alpha$ based on each user’s interactions, distilling user preferences to a user-personalized prompt in a compositional manner. Notably, our model can capture the complexities of a personalized user profile while maintaining a fixed number of prompts. Since the user-personalized prompt focuses on the most important parts of the user embedding, our method is similar high dimension clustering, meaning that users with similar preferences will have similar user-personalized prompts. We show how to obtain the user-personalized prompt in Figure 1.
4.4 Orthogonal Initialization
To reduce interference in knowledge between prompt components, we can initialize $P$, $K$, and $A$ orthogonally using QR Decomposition. Namely, let $Z\in\mathbb{R}^{R\times C}$ be a real matrix with $R$ rows and $C$ columns. $Z$ can be broken into an orthogonal matrix $Q$ and upper matrix $R$. The columns of $Q$ will thus be a set of $C$ orthogonal unit vectors which we can use as the initial vectors. We can repeat this initialization process for matrices $P$, $K$, and $A$.
4.5 Training and Generative Recommendation
Using the task-alternated training strategy proposed by Li et al. (2023), we alternate the task trained for each batch. This allows us to save time during training because the samples in the batch have the same data format and will therefore be a similar length to one another. Importantly, task-alternated training did not show any degradation in generative recommendation performance versus traditional random batching strategies Li et al. (2023). We combine user-personalized prompts, task-specific prompts, and discrete prompts as input during training. We also add whole-word embeddings to help the model identify which tokens belong to a particular ID. We show the whole mode architecture in Figure 2.
As all the generative recommendation tasks are posed as a sequence-to-sequence task, we use the Negative Log-Likelihood (NLL) loss to optimize the model parameters $\Theta$.
$\mathcal{L}_{\Theta}=\frac{1}{|\mathcal{D}|}\sum_{(X,Y)\in\mathcal{D}}\frac{1}{|Y|}\sum_{t=1}^{|Y|}-\log p(y_{t}|Y_{<t},X)$ (4)
where $\mathcal{D}$ is the training set that consists of input output pairs (X,Y) and $\log p(y_{t}|Y_{<t})$ is the probability of generating token $y_{t}$, given input sequence $X$ and the tokens previously generated.
In addtion, the model must output a sequence of text that forms an item ID (or a natural language text in the case of explanation generation). For this task, we choose beam search, which is commonly used for its effectiveness in sequence-to-sequence generation. Suppose we set the number of beams to $b$. At each time step, are $b$ candidate sequences. In the next step, any word in the vocabulary $V$ are appended to the end of the candidate sequences, resulting in $b\times V$ combinations. We can then select $b$ sequences that have the maximum log-likelihood. The LLM can continue this process until the candidate sequences reach a predefined maximum length. For sequential recommendation and top-$N$ recommendation, the $b$ candidate sequences form the recommendation list. For explanation, we select the sequence with the largest log-likelihood from the candidates.
5 Experimental Setup
5.1 Datasets
For experimentation, we use the Sports, Beauty, and Toys & Games Dataset from Amazon  [^1] https://jmcauley.ucsd.edu/data/amazon/. We reuse the preprocessing steps and the 8:1:1 train/validation/test splits created by Geng et al. (2022). Each sample consists of a user ID, an item ID, a text review, a rating, and a time stamp. We provide the dataset statistics in Table 1. For explanation generation, the Sentires toolkit Zhang et al. (2014) was used to extract ground-truth explanations for a given (user, item) pair. For sequential recommendation, the last item in a user’s history is part of the test set, the penultimate item is part of the validation set, and all other items are included in the train set. To prevent data leakage, the same training split used for sequential recommendation is used for top-n recommendation. For reproducibility, we provide all code used in the supplementary materials section of our submission and will release the code publicly upon publication.
Dataset Sports Beauty Toys
#Users 35,598 22,363 19,412
#Items 18,357 12,101 11,924
#Reviews 296,337 198,502 167,597
#Sparsity (%) 0.0453 0.0734 0.0724
Table 1: Statistics of the datasets.
5.2 Baselines
For sequential recommendation, we compare our model with the following baselines: Caser Tang and Wang (2018), HGN Ma et al. (2019), GRU4Rec Jannach and Ludewig (2017), BERT4Rec Sun et al. (2019), FDSA Zhang et al. (2019), SASRec Kang and McAuley (2018), P5 Geng et al. (2022), VIP5 Geng et al. (2023), and POD Li et al. (2023).
For top-n recommendation, we compare our method with: MF Koren et al. (2009), MLP Cheng et al. (2016), P5 Geng et al. (2022), VIP5 Geng et al. (2023), and POD Li et al. (2023).
Finally, for explanation generation, we use the following baselines: Att2Seq Dong et al. (2017):, NRT Li et al. (2020), PETER Li et al. (2021), and POD Li et al. (2023). For a fair comparison, we exclude P5 and VIP5 because we are focused exclusively on models that only use user and item IDs rather than additional user or item metadata.
Table 2: Performance comparison on sequential recommendation.
Methods Sports Beauty Toys
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
Caser 0.0116 0.0072 0.0194 0.0097 0.0205 0.0131 0.0347 0.0176 0.0166 0.0107 0.0270 0.0141
HGN 0.0189 0.0120 0.0313 0.0159 0.0325 0.0206 0.0512 0.0266 0.0321 0.0221 0.0497 0.0277
GRU4Rec 0.0129 0.0086 0.0204 0.0110 0.0137 0.0099 0.0283 0.0137 0.0097 0.0059 0.0176 0.0084
BERT4Rec 0.0115 0.0075 0.0191 0.0099 0.0203 0.0124 0.0347 0.0170 0.0116 0.0071 0.0203 0.0099
FDSA 0.0182 0.0122 0.0288 0.0156 0.0267 0.0163 0.0407 0.0208 0.0228 0.0140 0.0381 0.0189
SASRec 0.0233 0.0154 0.0350 0.0192 0.0387 0.0249 0.0605 0.0318 0.0463 0.0306 0.0675 0.0374
S3-Rec 0.0251 0.0161 0.0364 0.0204 0.0387 0.0247 0.0607 0.0327 0.0443 0.0294 0.0640 0.0376
P5 0.0272 0.0169 0.0361 0.0198 0.0530 0.0370 0.0659 0.0421 0.0460 0.0567 0.0709 0.0587
VIP5 0.0412 0.0345 0.0475 0.0365 0.0556 0.0427 0.0677 0.0467 0.0662 0.0577 0.0749 0.0604
POD 0.0496 0.0396 0.0576 0.0419 0.0537 0.0395 0.0688 0.0443 0.0691 0.0599 0.0742 0.0610
PeaPOD 0.0505 0.0400 0.0611 0.0432 0.0588 0.0445 0.0738 0.0493 0.0692 0.0583 0.0787 0.0609
Table 3: Performance comparison on top-n recommendation.
Methods Sports Beauty Toys
HR@1 HR@5 NDCG@5 HR@10 NDCG@10 HR@1 HR@5 NDCG@5 HR@10 NDCG@10 HR@1 HR@5 NDCG@5 HR@10 NDCG@10
MF 0.0314 0.1404 0.0848 0.2563 0.1220 0.0311 0.1426 0.0857 0.2573 0.1224 0.0233 0.1066 0.0641 0.2003 0.0940
MLP 0.0351 0.1520 0.0927 0.2671 0.1296 0.0317 0.1392 0.0848 0.2542 0.1215 0.0252 0.1142 0.0688 0.2077 0.0988
P5 0.0567 0.1514 0.1049 0.2196 0.1269 0.0571 0.1566 0.1078 0.2317 0.1318 0.0451 0.1322 0.0889 0.2023 0.1114
VIP5 0.0699 0.1882 0.1304 0.2717 0.1572 0.0615 0.1655 0.1147 0.2407 0.1388 0.0433 0.1301 0.0875 0.2037 0.1110
POD 0.0895 0.2086 0.1506 0.2873 0.1756 0.0829 0.1926 0.1391 0.2670 0.1629 0.0567 0.1433 0.1009 0.2082 0.1215
PeaPOD 0.1212 0.2741 0.2007 0.3616 0.2286 0.1097 0.2441 0.1780 0.3260 0.2039 0.0728 0.1618 0.1190 0.2315 0.1411
Table 4: Performance comparison on explanation generation.
Methods Sports Beauty Toys
BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
Att2Seq 0.5305 12.2800 1.2107 9.1312 0.7889 12.6590 1.6820 9.7481 1.6238 13.2245 2.9942 10.7398
NRT 0.4793 11.0723 1.1304 7.6674 0.8295 12.7815 1.8543 9.9477 1.9084 13.5231 3.6708 11.1867
PETER 0.7112 12.8944 1.3283 9.8635 1.1541 14.8497 2.1413 11.4143 1.9861 14.2716 3.6718 11.7010
POD 1.0013 14.0168 2.0436 11.1236 1.0630 15.2517 1.5737 11.3283 2.3053 12.2889 3.8512 10.3923
PeaPOD 0.9148 11.2685 1.7870 8.8078 1.0050 13.4532 1.5863 10.3238 2.4295 13.5537 3.9567 11.3863
5.3 Evaluation Metrics
For sequential and top-n recommendation, we evaluate models using two commonly used metrics, normalized Discounted Cumulative Gain (nDCG) and Hit Ratio (HR). nDCG is a metric used to evaluate the quality of ranked lists by considering the position of relevant items, while HR measures the fraction of relevant items that appear in the top-n positions of a ranked list. In particular, we measure HR@{5,10} and nDCG@{5,10} for sequential recommendation and HR@{1,5,10} and nDCG@{5,10} for top-n recommendation. For explanation generation, we evaluate models using BLEU Papineni et al. (2002) and ROUGE Lin and Hovy (2003). BLEU evaluates machine translation quality using n-gram precision against reference translations, while ROUGE assesses text summarization and translation by comparing n-gram overlaps with reference texts. We report the F1 scores for BLEU-4, ROUGE-1, ROUGE-2, and ROUGE-L. For all metrics, a greater score indicates better performance. In all tables, the best performing model for each metric is in bold and the second best model is underlined.
5.4 Implementation Details
To ensure a fair comparison, we compare PeaPOD with POD Li et al. (2023), P5 Geng et al. (2022), and VIP5 Geng et al. (2023) using T5-small (60.5 M parameters) Raffel et al. (2020) as the pre-trained LLM. On average, each experiment took 6 hours to run. Since the main focus of our study is not on formulating discrete prompts, we reuse the same discrete prompts as Geng et al. (2022). We adapt task-alternated training, as proposed by Li et al. (2023), to speed up training time by alternating each task per batch. Notably, both VIP5 and P5 both mention that fine-tuning with the T5-small pre-trained checkpoint often outperforms fine-tuning with T5-base.
For all models, we train the model on an Nvidia RTX 3090 for 30 epochs, a batch size of 64, the AdamW optimizer Kingma and Ba (2015), and a learning rate of 5e-3. We save the best performing model on the validation set. The length of the task-specific prompts and prompt components are both set to 3. We set the number of prompt components to 20. For our hyperparameter search, we tested prompt lengths of {3, 5, 8, 10} and the number of prompts of {10, 20, 30, 100}.To obtain the user embedding, we train a probabalistic matrix factorization model Mnih and Salakhutdinov (2007), a variant of matrix factorization, for 100 iterations with an embedding size of 512, learning rate of 1e-3, and a lambda regularization of 1e-3. During inference time, the number of beams is set to 20 and the inference batch size is set to 32 for all tasks.
6 Results
6.1 Sequential Recommendation
In Table 2, we compare our models with our selected baseline models. Our results show that PeaPOD outperforms all baselines in the Sports and Beauty datasets and is either the first or second best model for all metrics in the Toys dataset. Notably, the only difference between our method and POD, which is the second best performing model, is the additional personalized-user prompt that is prepended to the input. Interestingly, even though the user embedding passed to the attention mechanism is trained on a top-n recommender model, the additional information still provides a performance boost to sequential recommendation. This performance gain may be attributed to the way the data was split. In particular, the sequential and top-n recommendation contains the same user-item interactions, meaning that information about the users can be shared to improve performance in both tasks.
6.2 Top-n Recommendation
In Table 3, we report the performance of the top-n recommendation task. We note that PeaPOD performs the best in all metrics for all datasets, with a significant improvement in performance. We note that matrix factorization, which is used to generate the user embedding, is trained for top-n recommendation. Thus, the boost in performance may be more significant when compared to the other two tasks because of the user embedding selected as the query function. Nevertheless, we also find that PeaPOD outperforms matrix factorization itself, meaning that combining traditional collaborative filtering methods with LLMs leads to increased performance.
Figure 3: Comparison between PeaPOD and an ablated version (PeaPOD-X) trained without the task-specific prompts on our three tasks.
6.3 Explanation Generation
In Table 4, we report the performance of all models on the explanation generation task. Interestingly PeaPOD performs particularly well on both Sports and Toys datasets, but performs significantly worse in Beauty when compared to the top baseline models. Nevertheless, PeaPOD ranks first or second best on six out of the twelve evaluations, meaning that it is still competitive in explanation generation.
Upon manual inspection across all models, we notice that there are a significant amount of duplicates across explanations, meaning that the same explanations are generated regardless of the (user, item) pair. Thus, we posit that limiting the explanation generation to exclusively user and item IDs hurts performance. Similar to how performance for sequential and top-n recommendation was improved by providing collaborative user embeddings, the performance of item explanations can be improved by providing additional item metadata to the model. Previous research Geng et al. (2023); Tan et al. (2024); Rajput et al. (2023) has already explored how discrete item IDs can be transformed to provide more informative tokens to the model. However, we leave the additional experiments of combining enhanced item IDs with PeaPOD to future work.
6.4 Training without Task-Specific Prompts
To understand the effectiveness of user-personalized prompts versus globally shared prompts, we fine-tune the model with the same architecture and hyperparameters, but omit the task-specific prompts. We refer to the ablated model as ‘PeaPOD-X’.
As shown in Figure 3, we find that in all three datasets, the performance of the sequential recommendation marginally decreases, but the top-n recommendation performance improves. For explanation generation, the original model performs better for the Beauty dataset, but worse for Sports and Toys. We attribute the boost in top-n recommendation performance to the user embedding used for the query function. In particular, matrix factorization is a top-n recommendation model, meaning that the model is most likely to excel in this task. However, the lack of a task-specific prompt for sequential recommendation hinders the performance on this task. Nevertheless, we observe for the sequential recommendation task, the ablated model is still better than solely using the discrete prompt and has similar performance to POD. Thus, we show the effectiveness of a user-personalized prompt that utilizes preferences shared between similar users to improve recommendation performance. We report all metrics for this experiment in Appendix B.
7 Conclusion
In this paper, we propose PeaPOD, a novel recommender model that dynamically weights a set of decomposed prompt components to distill user preferences in a compositional manner. We maintain a many-to-many mapping between user preferences and continuous prompts to enable knowledge sharing among users with similar preferences. We utilize an attention-based mechanism to weight prompt components based on the user’s interaction history, thus creating a user-personalized prompt that distills both the user’s and similar users’ preferences into a learnable soft prompt. In our experiments, we show that our user-personalized prompt is effective in increasing the performance for three common recommendation tasks. Our model shows the effectiveness of preference distillation to improve personalized generative recommendation.
8 Limitations and Ethical Concerns
One limitation to our approach is that we limit our study to exclusively use user and item IDs. Thus, we do not incorporate any additional metadata such as user profiles or item descriptions. We find that this may hinder performance on explanation generation, which relies on having a strong understanding of both the user and item to provide a personalized response. We predict that providing additional metadata to the model can improve performance, particularly if it is given in natural language text as the LLM is already pre-trained on such data. We plan to extend PeaPOD to utilize this information in future work.
In addition, our model takes a two-step approach to generate user-personalized prompts. Namely, the user embedding is generated using matrix factorization, which is then passed to the model as the query function. However, the models are not connected end-to-end, which means that the user embedding cannot be continuously updated during training. For future work, we plan to combine these two models into a single end-to-end framework.
Another concern is the issue of fairness in generative recommender systems. For example, prior research has shown that LLM-based recommenders skew towards popular items, which causes a disparity in recommendations and can lead to a bias-amplification loop Yoon et al. (2024). In addition, past research has investigated the issue of fairness in the LLM itself Salutari et al. (2023). Thus, careful consideration must be made when deploying a generative recommender model to mitigate unfair recommendations.
Appendix A Baselines
For each of our three tasks, we compare our model with the following baselines.
A.1 Sequential Recommendation
Sequential recommendation predicts a user’s next preferred items based on their previous interactions, which are given in historical order.
  • Caser: ConvolutionAl Sequence Embedding Recommendation Tang and Wang (2018) uses a convolutional neural network to learn the sequential pattern of the user interaction history.
  • HGN Ma et al. (2019): Hierarchical Gating Network uses two gating modules to model a user’s short and long term interests.
  • GRU4Rec Jannach and Ludewig (2017): Gated Recurrent Unit (GRU) for Recommendation uses a GRU to process a user’s recommender session.
  • BERT4Rec Sun et al. (2019): Fine-tunes a bidirectional encoder representations from transformers model for sequential recommendation.
  • FDSA Zhang et al. (2019): Feature-level Deeper Self-Attention network. Integrates item features as feature sequences, which are then combined with item sequences to generate recommendations.
  • SASRec Kang and McAuley (2018): Self-Attentive based Sequential Recommendation model. Uses self-attention mechanisms for short-term semantics and Recurrent Neural Networks for long-term semantics.
  • $\mathbf{S^{3}}$-Rec Zhou et al. (2020): Self-Supervised learning for Sequential Recommendation. Uses four self-supervised learning objectives to capture the correlations between items in sequential data.
[…]
  • VIP5 Geng et al. (2023): An extension of P5 that trains on multimodal data.
  • POD Li et al. (2023): An extension of P5 that uses a task-specific, distilled prompt as part of the input.
A.2 Top-n Recommendation
Top-n recommendation predicts a list of the top-n items based on the user’s interaction history.
  • MF Koren et al. (2009): Matrix Factorization is a collaborative filtering technique that decomposes a user-item interaction matrix into latent user and item embeddings.
  • MLP Cheng et al. (2016): Multi-Layer Perceptron is a neural network trained to learn captures non-linear interactions between users and items.
  • P5 Geng et al. (2022): See definition in Section A.1.
  • VIP5 Geng et al. (2023): See definition in Section A.1.
  • POD Li et al. (2023): See definition in Section A.1.
A.3 Explanation Generation
Given a user ID and item ID, the model must return a natural language explanation describing why the user might like/dislike the target item. Notably, we exclude P5 and VIP5 because we are exclusively focused on models that exclusively use only user and item IDs rather than item titles or other item metadata.
  • Att2Seq Dong et al. (2017): an attention based model that learns to generate product reviews from item attributes.
  • NRT Li et al. (2020): Neural Ratings and Tips generation simultaneously returns a rating and a tip based on a neural network.
  • PETER Li et al. (2021): a transformer based model that performs rating prediction and explanation generation.
Appendix B Training without Task-Specific Prompts
We report the original model including task-specific prompts (denoted as PeaPOD) versus training the model without them (denoted as PeaPOD-X). All inputs and hyperparameters are otherwise the same.
Table 5: Performance comparison on sequential recommendation.
Methods Sports Beauty Toys
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
PeaPOD 0.0505 0.0400 0.0611 0.0432 0.0588 0.0445 0.0738 0.0493 0.0692 0.0583 0.0787 0.0609
PeaPOD-X 0.0480 0.0381 0.0581 0.0412 0.0565 0.0402 0.0732 0.0456 0.0705 0.0557 0.0835 0.0594
Table 6: Performance comparison on top-n recommendation.
Methods Sports Beauty Toys
HR@1 HR@5 NDCG@5 HR@10 NDCG@10 HR@1 HR@5 NDCG@5 HR@10 NDCG@10 HR@1 HR@5 NDCG@5 HR@10 NDCG@10
PeaPOD 0.1212 0.2741 0.2007 0.3616 0.2286 0.1097 0.2441 0.1780 0.3260 0.2039 0.0728 0.1618 0.1190 0.2315 0.1411
PeaPOD-X 0.1232 0.2744 0.2017 0.3662 0.2309 0.1130 0.2480 0.1820 0.3304 0.2079 0.0790 0.1795 0.1308 0.2513 0.1536
Table 7: Performance comparison on explanation generation.
Methods Sports Beauty Toys
BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
PeaPOD 0.9148 11.2685 1.7870 8.8078 1.0050 13.4532 1.5863 10.3238 2.4295 13.5537 3.9567 11.3863
PeaPOD-X 0.9561 12.1112 1.7613 9.4255 0.9598 11.3673 1.7317 8.8236 2.5306 15.3696 3.8183 12.6750