### v2

# PeaPOD: Personalized Prompt Distillation for Generative Recommendation

Abstract.
Large Language Models (LLMs) have become a powerful foundation for generative recommender systems, offering a unified approach to diverse recommendation tasks. However, existing multi-task LLM recommendation methods often rely on discrete ID-based prompts or task-specific soft prompts, which overlook the valuable collaborative signals shared among users with similar interests. To address this limitation, this paper presents a compositional framework that integrates a user’s individual preferences with collective preferences from similar users to build personalized soft prompts. Specifically, an attention-based mechanism fuses the embeddings of users with similar interests, creating a richer representation that captures multiple facets of user preferences. This design dynamically emphasizes shared interests while maintaining each user’s distinctive preferences. Experiments on three real-world datasets demonstrate the effectiveness of the proposed approach across sequential recommendation, top-n recommendation, and explanation generation tasks, underscoring the advantages of incorporating collaborative signals through an attention-based compositional strategy.
Recommender Systems, Personalization, Generative Recommendation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
1. Introduction
Recommender systems have emerged as pivotal tools in guiding users through vast amounts of items and tailoring personalized experiences across various tasks, including sequential recommendation, top-n recommendation, and explanation generation (Ge et al., 2022). Recently, recommender systems trained using large language models (LLMs), known as generative recommender systems, have unlocked the ability to perform multiple tasks with a single model. This unified design allows LLM-based recommender systems to perform well in various downstream tasks, including in a zero-shot or few-shot setting (Geng et al., 2022, 2023).
P5  (Geng et al., 2022) formulates recommendation as a text-to-text sequence task and trains the model to perform generative recommendations based on discrete prompt templates. Li et al. (2023) further proposed Prompt Distillation to address the disconnect between the IDs (including user and item) used in the discrete template and their original natural language meaning. In particular, they employ a continuous prompt to capture the information of the discrete prompt in a more flexible manner. However, existing work maintains prompts that are reused for all users. This globally-shared prompt design is in conflict with the fact that users’ personalized preferences greatly differ across the entire user population (Wu et al., 2022). In the real world, users might have similar or distinct preferences behind their interactions.
One intuitive solution for personalizing generative recommender systems is to create a personalized prompt for each user based solely on their individual interactions.  Tan et al. (2024b) have shown that fine-tuning a unique parameter-efficient fine-tuning (PEFT) module on each user’s interaction history can effectively personalize LLMs for personalized text generation and classification. However, this method is is reliant on the user having extensive historical data. Thus, fine-tuning a unique prompt per user is ineffective when the user’s interaction history is too sparse to enable effective personalization. Additionally, prior research in recommender systems has shown that sharing collaborative information among users with similar interests can improve performance in warm-start settings (Koren et al., 2009; Mnih and Salakhutdinov, 2007; Cheng et al., 2016).
Figure 1. Comparison of soft prompt architectures for multi-task generative recommendation. Unlike previous methods, PeaPOD utilizes the query-key-value mechanism from attention to generate a collaborative user prompt that captures group-level preferences based on the target user’s individual preferences.
In this work, we propose PErsonAlized PrOmpt Distillation (PeaPOD), a flexible framework that fuses individual and group-level signals into a unified collaborative user prompt for recommendation. To construct this system, we first utilize standard collaborative filtering techniques to generate initial user embeddings. These embeddings are then refined using a compositional, attention-based strategy that captures a contextualized representation of user-to-user relationships. Specifically, the user embedding of target user $u$ serves as the query vector in a multi-head attention module, while the embeddings of the top-n most similar users, identified via cosine similarity, act as both key and value vectors. By dynamically weighting shared preferences among similar users, our model generates a collaborative user soft prompt that effectively captures and aligns with the target user’s preferences. Moreover, we implement multiple attention heads to allow the model to capture different dimensions of similarity, resulting in a richer representation of the target user. As a result, PeaPOD provides a personalized soft prompt that can be used effectively across multiple generative recommendation tasks. Across three Amazon datasets, PeaPOD is able to achieve state-of-the-art performance on three recommendation tasks. We provide all data and code used at: https://anonymous.4open.science/r/peapod-3FF7/.
Our key contributions are summarized below:
  1. (1)
    We propose PErsonAlized PrOmpt Distillation (PeaPOD), a novel compositional strategy that distills collaborative user preferences into a personalized soft prompt for LLM-based recommendation.
  2. (2)
    We leverage the query-key-value mechanism from multi-head attention to capture diverse user-to-user relationships and construct a contextualized user representation that integrates both individual- and group-level information.
  3. (3)
    We conduct extensive experiments on sequential recommendation, top-n recommendation, and explanation generation tasks. Comprehensive ablation studies and results further validate the effectiveness and superiority of our proposed PeaPOD framework.
2. Related Works
2.1. LLMs for Recommendation
In recent years, researchers have explored the effectiveness of using LLMs for recommendation. The two main methods of LLMs for recommender systems are (1) using the LLM’s pre-trained knowledge to recommend items in a zero-shot or few-shot setting and (2) fine-tuning the model for a given domain.
One advantage of LLMs is that they are pre-trained with a large amount of data from various domains. Thus,  Sanner et al. (2023) crowdsourced a dataset of natural language summaries of user preferences and showed that prompting a language model with these user summaries yielded competitive performance with traditional baselines in a near cold-start setting.  Ramos et al. (2024) showed that a generative recommendation system trained solely on natural language user profiles are also competitive with traditional recommender systems in a warm-start setting, with the added benefit of transparency and scrutability. In addition, the zero-shot prompting of GPT-4 has also beaten SOTA models in a conversational recommender setting (He et al., 2023). Although prompting LLMs can lead to strong performance in recommender tasks, they are reliant on better content knowledge rather than collaborative knowledge (He et al., 2023; Wu et al., 2024). Thus, if the item metadata is not already learned in the pre-trained knowledge base or if the target dataset users interact differently than the users that the LLM was originally trained on, the performance of the recommender system will suffer.
To mitigate these issues, the model can be further fine-tuned on target datasets in order to train the model based on the given user population and add or update its knowledge on the item dataset. P5 (Geng et al., 2022) is the first work that trains a T5-based (Raffel et al., 2020) foundational recommender model to unify various recommendation tasks in a shared framework. VIP5 (Geng et al., 2023) extends P5 to a multi-modal foundational recommender system. POD (Li et al., 2023) builds upon the P5 model and appends a task-specific, continuous prompt to the input to improve performance. Although these models pass in the user ID as inputs, they still do not leverage the collaborative information contained from similar users. Zhu et al. (2024) expands the tokenizer vocabulary to add user/item IDs as tokens and learn collaborative information, but this technique is infeasible as the number of users/items grows. In this work, we focus on developing a foundational generative recommender model that can efficiently encode individual and collaborative preferences for multiple downstream tasks.
2.2. Structural Prompt
One issue with generative approaches to recommendation is how to effectively encode user-item interactions to use with LLMs. Researchers have proposed numerous methods to bridge the gap between IDs and natural language. One option is to add the user’s name and item title to the input prompt. However, this data may be unavailable or non-unique, making it difficult to personalize recommendations. Another option is to directly insert the user and item IDs to the discrete prompt template (Geng et al., 2022). However, LLMs are trained on natural language text rather than IDs, which limits the effectiveness of this approach (Li et al., 2023).
To tackle these issues, many researchers have explored how to effectively encode item IDs with LLMs.  Rajput et al. (2023) proposed using RQ-VAE with item metadata to generate semantically meaningful tuples of tokens to serve as the semantic ID for each item. GPTRec uses a novel SVD tokenization method to generate quantized item embeddings based on user-item interactions (Petrov and Macdonald, 2023).  Tan et al. (2024a) proposed training a textual ID generator alongside the LLM-based recommender to convert item IDs to meaningful text descriptions. While these methods are effective, they do not address the need for encoding user preferences as input to the model.
To incorporate user preferences into the LLM-based recommender, Zhang et al. (2024) proposes a mapping technique to project user embeddings learned from collaborative filtering into a soft prompt. In this work, we focus on developing an attention-based compositional method to better model user-to-user relationships. Previous work in continual learning for image classification has shown that an attention-based mechanism can be used to effectively share information between similar classes of images to improve performance (Smith et al., 2023). In the context of generative recommendation, multi-head attention (Vaswani et al., 2017) is used to capture both individual and collaborative preferences across diverse feature spaces.
3. Preliminary
3.1. Recommendation Tasks
Following existing works (Geng et al., 2022; Li et al., 2023), we focus on three important recommendation tasks: (1) Sequential Recommendation, (2) Top-n Recommendation, and (3) Explanation Generation. Let $u\in U$ denote a user in the user set and $i\in I$ denote an item in the item set across all three tasks. We aim to train a unified model to complete the following tasks:
  • Sequential Recommendation: given a user $u$ and a list of historical interactions in chronological order $i_{1},i_{2},...i_{n-1}$, the task is to predict the next item $i_{n}$ that $u$ will interact with.
  • Top-n Recommendation: given a pool of randomly sampled items from $I$, the model must select the top $N$ items that the user wants to interact with.
  • Explanation Generation: given a user $u$ and item $i$, the model must generate an explanation $E_{u,i}$ that explains why $u$ likes/dislikes $i$.
Importantly, the scope of our study focuses only on user and item IDs for recommendation and excludes additional information such as item metadata or user profiles because this information may be sensitive or unavailable.
Table 1. Notation Table
Symbol Description
$u$ Target user
$i$ Item
$E_{u,i}$ Explanation why $u$ likes/dislikes $i$
$\boldsymbol{F}$ User-Item feedback matrix
$d_{u}$ Dimension of user embedding
$S_{u}$ Top-n most similar users to $u$
$d_{m}$ Dimension of model
$\mathbf{Q}$, $\mathbf{K}$, $\mathbf{V}$ Query, key, and value matrices
$\mathbf{W}_{h}^{Q}$, $\mathbf{W}_{h}^{K}$, $\mathbf{W}_{h}^{V}$ Linear projection weights for head $h$
$\mathbf{z}$ User’s summary vector
$\mathbf{p}$ Collaborative user prompt
$d_{p}$ Dimension of collaborative user prompt
$\boldsymbol{head}_{h}$ Output of attention for head $h$
$d_{k}$ Dimension of attention head
$\mathbf{W}_{l}$ Linear layer weight for feedforward network
$\mathbf{b}$ Bias vector
$H$ Number of attention heads
3.2. Prompt Distillation
In traditional recommender systems, user and item IDs are commonly used to identify unique users and items. However, since these IDs are unique to the dataset and are not natural language text, it is necessary to perform additional fine-tuning to allow the LLM to effectively understand these IDs. The most common methodology is to train the model with discrete prompt templates. For example, in the case of explanation generation, the discrete prompt could be “Explain why user_1234 enjoys item_5678”. When using IDs with discrete prompt templates, recent research proposes adding an additional embedding space, called whole-word embeddings, so that the model can recognize which tokens belong to the appropriate ID (Geng et al., 2022). That is, each token in the ID utilizes the same whole-word embedding and all other tokens are tokenized as normal.
Although discrete prompts can lead to good performance, they are inflexible and require manually crafted prompts. Thus, the performance of the model may depend strongly on how well the instructions were written. On the other hand, continuous prompts are dynamically adjusted and fine-tuned through techniques like soft prompt tuning, which optimizes prompts by learning continuous embeddings. This approach allows the model to learn regardless of the quality of the prompt template. Li et al. (2023) introduced Prompt Distillation, which distills discrete prompts into continuous prompt vectors. The authors used prompt distillation to incorporate task-specific, continuous prompts for each recommendation task. These prompts are shared globally across all users to improve the model’s performance. Since this method adds a shared continuous prompt per recommendation task, we will refer to these continuous prompts as task-specific prompts. Although task-specific prompts were shown to improve performance in our recommendation tasks, they cannot express personalization because each user utilizes the same set of prompts. Thus, it is imperative to develop a new method to distill each user’s preferences into a personalized soft prompt that caters to each user’s individual needs, while incorporating relevant preferences from similar users.
4. Collaborative User Prompt
Figure 2. Full PeaPOD model architecture. We first initialize a set of user embeddings using probabilistic matrix factorization. We then apply multi-head attention, where the user embedding of target user $u$ is passed to the query vector and the embeddings of the top-n most similar users $S_{u}$ are passed to the key and value vectors. The final output is the collaborative user prompt $\boldsymbol{p}_{mha}$, which dynamically captures collaborative knowledge from $S_{u}$ based on $u$’s personalized preferences.
The goal of the collaborative user prompt is to distill user preferences into a soft prompt to represent the user’s personal interests while also leveraging collaborative knowledge from users with similar preferences. To accomplish this, we implement multi-head attention (Vaswani et al., 2017) to create a contextualized user representation that combines individual and group-level information. Attention dynamically weights the influence of each similar user, producing a richer and more robust representation. Furthermore, the utilization of multiple heads allows the model to capture different aspects of similarity, resulting in a more comprehensive and diverse understanding of the target user. By applying multi-head attention, we are able to generate a personalized soft prompt that can be effectively used and shared by the LLM across multiple recommender tasks.
4.1. User Embedding Representation
To generate the collaborative user prompt, we must first select an effective input to the multi-head-attention module. In our case, we choose to utilize a user embedding generated from probabilistic matrix factorization (Mnih and Salakhutdinov, 2007) to generate embeddings for each user. In particular, given a user-item feedback matrix $\boldsymbol{F}\in\mathbb{R}^{|\boldsymbol{U}|\times|\boldsymbol{I}|}$, the model learns a user matrix $\boldsymbol{U}\in\mathbb{R}^{|\boldsymbol{U}|\times d_{u}}$ and item matrix $\boldsymbol{I}\in\mathbb{R}^{|\boldsymbol{I}|\times d_{u}}$, where $d_{u}$ is the selected dimension size of the user embedding and $|\boldsymbol{U}|$ and $|\boldsymbol{I}|$ are the number of users and items, respectively. $\boldsymbol{u}\in\boldsymbol{U}$ represents a user embedding that encodes a user’s individual preferences. By construction, users with similar preferences end up closer in this embedding space (e.g. higher cosine similarity). These pretrained user embeddings serve as a strong initialization for personalization because they already capture collaborative signals.
4.2. Attending User Embeddings
In our framework, attention is utilized to generate a collaborative user prompt $\boldsymbol{p}$ by leveraging the embeddings of user $u$ (e.g. $\boldsymbol{u}$) and $\boldsymbol{S}_{u}\in\mathbb{R}^{n\times d_{u}}$, which is defined as the top-n most similar user embeddings to $\boldsymbol{u}$ by cosine similarity. We attend over these $n$ neighboring user embeddings to capture their collective preferences, conditioned on $u$. We first project $\boldsymbol{u}$ and each neighbor’s embedding $\boldsymbol{s}_{i}\in\boldsymbol{S}_{u}$ into the query $\boldsymbol{Q}$, key $\boldsymbol{K}$, and value $\boldsymbol{V}$ subspaces.
(1) $\boldsymbol{Q}$ $=\boldsymbol{W}_{Q}u+\boldsymbol{b}_{Q}\in\mathbb{R}^{d_{m}}$
(2) $\boldsymbol{K_{i}}$ $=\boldsymbol{W}_{K}s_{i}+\boldsymbol{b}_{K}\in\mathbb{R}^{d_{m}}$
(3) $\boldsymbol{V}_{i}$ $=\boldsymbol{W}_{V}s_{i}+\boldsymbol{b}_{V}\in\mathbb{R}^{d_{m}}$
where $d_{m}$ is the dimensionality of the base LLM. Aggregating similar users, we define the key and value spaces as:
(4) $\mathbf{K}=\begin{bmatrix}\mathbf{K}_{1}\\ \mathbf{K}_{2}\\ \vdots\\ \mathbf{K}_{n}\end{bmatrix}\in\mathbb{R}^{n\times d_{m}},\quad\mathbf{V}=\begin{bmatrix}\mathbf{V}_{1}\\ \mathbf{V}_{2}\\ \vdots\\ \mathbf{V}_{n}\end{bmatrix}\in\mathbb{R}^{n\times d_{m}}$
$\boldsymbol{Q}$ is the representation of the target user $u$ that we condition group-level information on. $\boldsymbol{K}$ encodes $\boldsymbol{S}_{u}$ in the key space, enabling the model to learn how relevant each similar user is to $u$. $\boldsymbol{V}$ carries the content from $\boldsymbol{S}_{u}$ to be aggregated by attention. We use scaled dot-product attention to compute an attention vector and apply it to $\boldsymbol{V}$:
(5) $\boldsymbol{z}=\text{Attention}(\boldsymbol{Q,K,V})=\text{softmax}\left(\frac{\boldsymbol{QK}^{\top}}{\sqrt{d_{m}}}\right)\boldsymbol{V}\in\mathbb{R}^{d_{m}},$
This results in an attention-weighted summary vector $\boldsymbol{z}$. In the context of our selected inputs, This allows us to dynamically aggregate the most important information from similar users with regards to $u$, further enhancing the collaborative information learned from the initialized user embeddings. Finally, we apply a linear layer to $\boldsymbol{z}$ to obtain the collaborative user prompt $\boldsymbol{p}$.
(6) $\boldsymbol{p}$ $=\boldsymbol{W}_{l}\boldsymbol{z}+\boldsymbol{b}_{l}\in\mathbb{R}^{d_{p}}$
where $d_{p}$ is the selected prompt length for the collaborative user prompt. By attending to the top-n most similar users, we dynamically incorporate relevant collaborative signals into $\boldsymbol{p}$ based on $u$. This compositional strategy combines group-level information, conditioned on $u$’s preferences, to distill collaborative preferences into a personalized soft prompt.
4.3. Multi-Head Attention
The collaborative user prompt described in Section 4.2 can be extended from a single-head implementation to a multi-head attention (Vaswani et al., 2017). Multi-head attention enables the model to attend to different subspaces of the embeddings simultaneously, allowing the model to capture richer relationships and patterns across the top-$n$ similar users.
Formally, instead of using a single set of linear projections for $\boldsymbol{Q}$, $\boldsymbol{K}$, and $\boldsymbol{V}$, we learn multiple sets of projections, one for each head $h\in\{1,\dots,H\}$:
(7) $\textbf{head}_{h}$ $=\text{Attention}\bigl{(}\boldsymbol{QW}_{h}^{Q},\;\boldsymbol{KW}_{h}^{K},\;\boldsymbol{VW}_{h}^{V}\bigr{)}\in\mathbb{R}^{d_{k}},$
where $\mathbb{R}^{d_{k}}=\mathbb{R}^{d_{m}/H}$ is the dimension of each attention head. After concatenating all heads, the final dimension remains $d_{m}$. The outputs of each head are then concatenated and passed through a linear layer:
(8) $\boldsymbol{z}_{mha}=\text{Concat}\bigl{(}\textbf{head}_{1},\dots,\textbf{head}_{H}\bigr{)}\in\mathbb{R}^{d_{m}}\$
(9) $\boldsymbol{p}_{mha}=\boldsymbol{W}_{l}\boldsymbol{z}_{mha}+\boldsymbol{b}_{l}\in\mathbb{R}^{d_{p}}$
In the context of the collaborative user prompt, multi-head attention is able to capture multiple, diverse user-to-user relationships between $u$ and $S_{u}$. Specifically, each head can learn to focus on different aspects of the user embeddings. For example, one head might focus on users who share highly overlapping item preferences, while another might capture niche or temporal similarities. By combining multiple heads, we integrate these different viewpoints into a more comprehensive collaborative user representation. Because multi-head attention allows us to capture more nuanced similarities among neighbors, it potentially leads to better personalization than single-head attention.
4.4. Training and Generative Recommendation
Using the task-alternated training strategy proposed by Li et al. (2023), we alternate the task trained for each batch. This allows us to save time during training because the samples in the batch have the same data format and will therefore be a similar length to one another. Importantly, task-alternated training did not show any degradation in generative recommendation performance versus traditional random batching strategies (Li et al., 2023). We combine collaborative user prompts, task-specific prompts, and discrete prompts as input during training. We also add whole-word embeddings to help the model identify which tokens belong to a particular ID (Geng et al., 2022). We show the whole model architecture in Figure 2.
As all the generative recommendation tasks are posed as a sequence-to-sequence task, we use the Negative Log-Likelihood (NLL) loss to optimize the model parameters $\Theta$.
(10) $\mathcal{L}_{\Theta}=\frac{1}{|\mathcal{D}|}\sum_{(X,Y)\in\mathcal{D}}\frac{1}{|Y|}\sum_{t=1}^{|Y|}-\log p(y_{t}|Y_{<t},X)$
where $\mathcal{D}$ is the training set that consists of input output pairs (X,Y) and $\log p(y_{t}|Y_{<t})$ is the probability of generating token $y_{t}$, given input sequence $X$ and the tokens previously generated.
In addition, the model must output a sequence of text that forms an item ID (or natural language text in the case of explanation generation). For this task, we choose beam search, which is commonly used for its effectiveness in sequence-to-sequence generation. Suppose we set the number of beams to $b$. At each time step, are $b$ candidate sequences. In the next step, any word in the vocabulary $\boldsymbol{V}$ are appended to the end of the candidate sequences, resulting in $b\times\boldsymbol{V}$ combinations. We can then select $b$ sequences that have the maximum log-likelihood. The LLM can continue this process until the candidate sequences reach a predefined maximum length. For sequential recommendation and top-n recommendation, the $b$ candidate sequences form the recommendation list. For explanation, we select the sequence with the largest log-likelihood from the candidates.
5. Experimental Setup
The following research questions will guide the remainder of the paper:
  • RQ1: Does our proposed PeaPOD model outperform existing state-of-the-art models on different tasks?
  • RQ2: Does using a different method to obtain user embeddings significantly change the recommendation performance?
  • RQ3: What is the effectiveness of adding a collaborative user prompt as input to the model?
  • RQ4: How does model size impact performance?
  • RQ5: How useful is multi-head attention vs. alternative implementations such as single-head attention or multi-layer perceptron?
5.1. Datasets
To answer the research questions, we use the Sports, Beauty, and Toys & Games Dataset from Amazon (Ni et al., 2019). We reuse the preprocessing steps and the 8:1:1 train/validation/test splits created by Geng et al. (2022). Each sample consists of a user ID, an item ID, a text review, a rating, and a time stamp. We provide the dataset statistics in Table 2. For explanation generation, the Sentires toolkit (Zhang et al., 2014) was used to extract ground-truth explanations for a given (user, item) pair. For sequential recommendation, the last item in a user’s history is part of the test set, the penultimate item is part of the validation set, and all other items are included in the train set. To prevent data leakage, the same training split used for sequential recommendation is used for top-n recommendation.
Dataset Sports Beauty Toys
#Users 35,598 22,363 19,412
#Items 18,357 12,101 11,924
#Reviews 296,337 198,502 167,597
#Sparsity (%) 0.0453 0.0734 0.0724
Table 2. Statistics of the datasets.
5.2. Baselines
We report the following baselines for each of the three recommendation tasks:
  • For sequential recommendation, we compare with the following baselines: Caser (Tang and Wang, 2018), HGN (Ma et al., 2019), GRU4Rec (Jannach and Ludewig, 2017),
    BERT4Rec (Sun et al., 2019), FDSA (Zhang et al., 2019), SASRec (Kang and McAuley, 2018), P5 (Geng et al., 2022), VIP5 (Geng et al., 2023), POD (Li et al., 2023), IDGenRec (Tan et al., 2024a), and RDRec (Wang et al., 2024).
  • For top-n recommendation, we compare our method with: MF (Koren et al., 2009), MLP (Cheng et al., 2016), P5 (Geng et al., 2022), VIP5 (Geng et al., 2023), POD (Li et al., 2023), and RDRec (Wang et al., 2024).
  • For explanation generation, we use the following baselines: Att2Seq (Dong et al., 2017):, NRT (Li et al., 2020), PETER (Li et al., 2021), P5 (Geng et al., 2022), VIP5 (Geng et al., 2023) and POD (Li et al., 2023).
5.3. Evaluation Metrics
For sequential and top-n recommendation, we evaluate models using two commonly used metrics, normalized Discounted Cumulative Gain (nDCG) and Hit Ratio (HR). nDCG is a metric used to evaluate the quality of ranked lists by considering the position of relevant items, while HR measures the fraction of relevant items that appear in the top-n positions of a ranked list. In particular, we measure HR@{5,10} and nDCG@{5,10} for sequential recommendation and HR@{1,5,10} and nDCG@{5,10} for top-n recommendation. For explanation generation, we evaluate models using BLEU (Papineni et al., 2002) and ROUGE (Lin and Hovy, 2003). BLEU evaluates machine translation quality using n-gram precision against reference translations, while ROUGE assesses text summarization and translation by comparing n-gram overlaps with reference texts. We report the F1 scores for BLEU-4, ROUGE-1, ROUGE-2, and ROUGE-L. For all metrics, a greater score indicates better performance. In all tables, the best performing model for each metric is in bold and the second best model is underlined.
5.4. Implementation Details
To ensure a fair comparison, we compare PeaPOD with POD (Li et al., 2023), P5 (Geng et al., 2022), and VIP5 (Geng et al., 2023) using T5-small (60.5 M parameters) (Raffel et al., 2020) as the base LLM. On average, each experiment took 6 hours to run. Since the main focus of our study is not on formulating discrete prompts, we reuse the same discrete prompts as POD (Li et al., 2023). We adapt task-alternated training, as proposed by Li et al. (2023), to speed up training time by alternating each task per batch.
To initialize user embeddings for PeaPOD, we train a probabilistic matrix factorization model (Mnih and Salakhutdinov, 2007) for top-n recommendation. We set the hyperparameters to 100 iterations with a dimension size of 512, learning rate of 1e-3, and a lambda regularization of 1e-3. We will refer to this model as PeaPOD-PMF. We train PeaPOD-PMF on an Nvidia RTX 3090 for 30 epochs. We utilize the AdamW optimizer (Kingma and Ba, 2015), with a batch size of 64, learning rate of 5e-3, and length of the task-specific prompt set to 3. We save the best performing model on the validation set.
For the collaborative user prompt, we perform a hyperparameter search, testing prompt lengths of {3, 5, 8, 10}, number of heads of {1, 2, 4, 8, 16} and the number of similar users of {5, 10, 20, 30}. For Sports and Beauty, the optimal hyperparameters we found was 20 similar users and the length of the collaborative user prompt $|\boldsymbol{p}_{mha}|=3$. For Toys, the optimal hyperparameters we found was 10 similar users and $|\boldsymbol{p}_{mha}|=3$. For all dataset, $H=4$. During inference time, the number of beams is set to 20 and the inference batch size is set to 32 for all tasks. We include our implementation at: https://anonymous.4open.science/r/peapod-3FF7/.
6. Results
6.1. Recommendation Performance (RQ1)
6.1.1. Sequential Recommendation
In Table 3, we compare our models with the baseline models on the sequential recommendation task. Our results show that PeaPOD-PMF outperforms all baselines in the three datasets. In addition, the superior performance of our model shows that user embeddings generated from one recommendation task can act as a strong signal for other recommender tasks. Although the set of user embeddings is generated from PMF, a top-n recommendation model, PeaPOD-PMF outperforms traditional sequential recommendation models. Thus, PeaPOD can effectively utilize personalized information from PMF to perform recommendation tasks that PMF is not explicitly trained for.
Table 3. Performance comparison on sequential recommendation.
Methods Sports Beauty Toys
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
Caser 0.0116 0.0072 0.0194 0.0097 0.0205 0.0131 0.0347 0.0176 0.0166 0.0107 0.0270 0.0141
HGN 0.0189 0.0120 0.0313 0.0159 0.0325 0.0206 0.0512 0.0266 0.0321 0.0221 0.0497 0.0277
GRU4Rec 0.0129 0.0086 0.0204 0.0110 0.0137 0.0099 0.0283 0.0137 0.0097 0.0059 0.0176 0.0084
BERT4Rec 0.0115 0.0075 0.0191 0.0099 0.0203 0.0124 0.0347 0.0170 0.0116 0.0071 0.0203 0.0099
FDSA 0.0182 0.0122 0.0288 0.0156 0.0267 0.0163 0.0407 0.0208 0.0228 0.0140 0.0381 0.0189
SASRec 0.0233 0.0154 0.0350 0.0192 0.0387 0.0249 0.0605 0.0318 0.0463 0.0306 0.0675 0.0374
S3-Rec 0.0251 0.0161 0.0364 0.0204 0.0387 0.0247 0.0607 0.0327 0.0443 0.0294 0.0640 0.0376
P5 0.0272 0.0169 0.0361 0.0198 0.0530 0.0370 0.0659 0.0421 0.0460 0.0567 0.0709 0.0587
VIP5 0.0412 0.0345 0.0475 0.0365 0.0556 0.0427 0.0677 0.0467 0.0662 0.0577 0.0749 0.0604
IDGenRec 0.0429 0.0326 0.0574 0.0372 0.0618 0.0486 0.0814 0.0541 0.0655 0.0481 0.0870 0.0551
POD 0.0496 0.0396 0.0576 0.0419 0.0537 0.0395 0.0688 0.0443 0.0691 0.0599 0.0742 0.0610
RDRec 0.0505 0.0408 0.0596 0.0433 0.0601 0.0461 0.0743 0.0504 0.0723 0.0593 0.0802 0.0605
PeaPOD-PMF 0.0534 0.0424 0.0638 0.0456 0.0718 0.0527 0.0915 0.0591 0.0865 0.0708 0.1018 0.0750
Table 4. Performance comparison on top-n recommendation.
Methods Sports Beauty Toys
HR@1 HR@5 NDCG@5 HR@10 NDCG@10 HR@1 HR@5 NDCG@5 HR@10 NDCG@10 HR@1 HR@5 NDCG@5 HR@10 NDCG@10
MF 0.0314 0.1404 0.0848 0.2563 0.1220 0.0311 0.1426 0.0857 0.2573 0.1224 0.0233 0.1066 0.0641 0.2003 0.0940
MLP 0.0351 0.1520 0.0927 0.2671 0.1296 0.0317 0.1392 0.0848 0.2542 0.1215 0.0252 0.1142 0.0688 0.2077 0.0988
P5 0.0567 0.1514 0.1049 0.2196 0.1269 0.0571 0.1566 0.1078 0.2317 0.1318 0.0451 0.1322 0.0889 0.2023 0.1114
VIP5 0.0699 0.1882 0.1304 0.2717 0.1572 0.0615 0.1655 0.1147 0.2407 0.1388 0.0433 0.1301 0.0875 0.2037 0.1110
POD 0.0895 0.2086 0.1506 0.2873 0.1756 0.0829 0.1926 0.1391 0.2670 0.1629 0.0567 0.1433 0.1009 0.2082 0.1215
RDRec 0.1285 0.2747 0.2033 0.3683 0.2326 0.1203 0.2572 0.1902 0.3380 0.2160 0.0660 0.1655 0.1171 0.2375 0.1398
PeaPOD-PMF 0.1252 0.2708 0.2006 0.3596 0.2287 0.1292 0.2628 0.1978 0.3423 0.2228 0.0728 0.2126 0.1581 0.2872 0.1817
Table 5. Performance comparison on explanation generation.
Methods Sports Beauty Toys
BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
Att2Seq 0.5305 12.2800 1.2107 9.1312 0.7889 12.6590 1.6820 9.7481 1.6238 13.2245 2.9942 10.7398
NRT 0.4793 11.0723 1.1304 7.6674 0.8295 12.7815 1.8543 9.9477 1.9084 13.5231 3.6708 11.1867
PETER 0.7112 12.8944 1.3283 9.8635 1.1541 14.8497 2.1413 11.4143 1.9861 14.2716 3.6718 11.7010
P5 1.0447 14.9048 2.1297 11.1778 1.2237 17.6938 2.2489 12.8606 2.2892 15.4505 3.6974 12.1718
VIP5 1.0639 14.8628 2.1012 11.1059 1.2850 17.7492 2.3482 12.9170 2.3241 15.3006 3.6590 12.0421
POD 1.0013 14.0168 2.0436 11.1236 1.0630 15.2517 1.5737 11.3283 2.3053 12.2889 3.8512 10.3923
PeaPOD-PMF 1.0313 13.7253 1.8260 10.7078 1.0468 12.9759 1.8853 10.2740 2.5145 13.8132 4.0357 11.5514
6.1.2. Top-n Recommendation
In Table 4, we report the performance of our method and the baselines on the top-n recommendation task. We found that PeaPOD-PMF outperformed existing baselines on the Beauty and Toys datasets and placed second in Sports. Although PeaPOD-PMF was initialized on a variant of matrix factorization, PeaPOD-PMF showed significant gains over the matrix factorization baseline. Thus, our attention-based compositional strategy is able to further refine the collaborative information provided by the initial user embedding. We explore the effects of the selected method to initialize user embeddings in Section 6.2.
6.1.3. Explanation Generation
In Table 5, we report the performance of all models on the explanation generation task. We observe that PeaPOD-PMF performs competitively in Toys and Sports. In particular, PeaPOD-PMF achieved the best performance in BLEU-4 and ROUGE-2 in Toys. However, PeaPOD-PMF performs worse in Beauty. Importantly, the top-two models, P5 and VIP5, contained additional item metadata in the input that the other models did not have. In particular, P5 provides the item title in the prompt and VIP5 includes both the item title and the item photo. Thus, we posit that including item metadata in the input greatly contributes to the improvement of performance for explanation generation. Although the scope of our study focuses on scenarios where only user and item IDs are available to the model, PeaPOD can be easily extended to include these additional metadata as input. Integrating these complementary methods can lead to better overall performance, making it a compelling avenue for future research.
6.2. Sensitivity to User Embeddings (RQ2)
We train an ablated variant of PeaPOD, named PeaPOD-BiVAE, to understand the effect that initial user embeddings have on the PeaPOD architecture. We initialize the user embeddings of PeaPOD-BiVAE by training a Bilateral Variational Autoencoder (BiVAE) (Truong et al., 2021) for top-n recommendation. We train BiVAE for 100 epochs with an embedding size of 512, learning rate of 1e-3, and $\lambda$ of 1e-3. We report the results of both PeaPOD models in Figure 3.
Overall, we find that the method of initializing user embeddings affects how well the model performs on the different tasks. PeaPOD-BiVAE performed better on sequential recommendation for Sports, but PeaPOD-PMF performed better on all other metrics. Thus, the method used to obtain the initial user embeddings plays an important role in the overall recommendation performance.
It is important to note that the user embedding does not need to be generated from a traditional recommender system. The primary reason we chose to initialize user embeddings using collaborative filtering was that it provides key information about user-to-user interactions that the attention mechanism can further refine. However, any method that provides a strong signal about personalized and collaborative preferences can be used as input to PeaPOD. For future work, we plan to optimize the user embedding initialization and collaborative user prompt generation steps into a unified, end-to-end architecture so that the two components can be trained at the same time rather than splitting them into two distinct steps.
Figure 3. Comparison between PeaPOD-PMF and PeaPOD-BiVAE.
Figure 4. Comparison between PeaPOD-PMF and an ablated version, PeaPOD (no task), trained without task-specific prompts.
6.3. Effects of Collaborative User Prompt (RQ3)
To understand the effectiveness of collaborative user prompts versus globally shared prompts, we fine-tune PeaPOD-PMF with the same architecture and hyperparameters, but omit the task-specific prompts. We refer to the ablated model as PeaPOD-PMF (no task). We report our findings in Figure 4.
We find that for sequential and top-n recommendation, PeaPOD-PMF performs better on Beauty and Toys, whereas PeaPOD-PMF (no task) performs better in Sports. For explanation generation, the models perform similar to each other, with a slight edge for PeaPOD-PMF in Beauty and PeaPOD-PMF (no task) in Toys. Importantly, we find that PeaPOD-PMF (no task) shows better performance overall versus POD in sequential recommendation and top-n recommendation. Thus, the personalized information distilled into the collaborative user prompt encodes more relevant information for generative recommendation compared to the task-specific prompt. However, combining the two types of prompts leads to better performance overall. Since both models are fine-tuned using same LLM (T5-small) and discrete prompt templates, we can attribute the gain in PeaPOD’s performance to the effectiveness of a collaborative user prompt that utilizes preferences shared between similar users to improve recommendation performance.
6.4. Effects of Model Size (RQ4)
We investigate the effects of a larger base model size on the recommendation performance. In particular, we compare PeaPOD-PMF (small) and PeaPOD-PMF (base), which are models fine-tuned on T5-small (60.5M params) and T5-base (223M params), respectively. We train both models with the same hyperparameters and user embeddings. We report our findings in Tables 6, 7, and 8.
For sequential recommendation, we find that PeaPOD-PMF (base) outperforms PeaPOD-PMF (small) on the Sports dataset, but underperforms on the Beauty and Toys datasets. For top-n recommendation, PeaPOD-PMF (base) performs worse for the Beauty and Toys datasets, but is competitive with PeaPOD-PMF (small) in Sports. In particular, PeaPOD-PMF (base) saw the lowest drop in performance on the Toys dataset. For explanation generation, PeaPOD-PMF (base) performs better for Sports and PeaPOD-PMF (small) performs better for Beauty. Interestingly, we find that PeaPOD-PMF (base) overfits on the model much earlier than its small counterpart on all three datasets.
Overall, we find that a bigger model does not always equate to a better performance. In fact, PeaPOD-PMF (small) model performs better on the majority of metrics despite having fewer parameters.  Geng et al. (2022) found a similar finding, noting that the smaller P5 variant outperformed the larger P5 variant on numerous tasks. In addition, both our and Geng et al. (2022)’s experiments found that the larger model performs significantly better on the Sports dataset, which has significantly more data and a lower sparsity compared to Beauty and Toys datasets. These findings indicate that a larger model may be needed when there is more data, whereas a smaller model might be more appropriate when training on a smaller dataset with higher sparsity.
Table 6. Performance comparison on sequential recommendation between small and base models.
Methods Sports Beauty Toys
HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10 HR@5 NDCG@5 HR@10 NDCG@10
PeaPOD-PMF (small) 0.0534 0.0424 0.0638 0.0456 0.0718 0.0527 0.0915 0.0591 0.0865 0.0708 0.1018 0.0750
PeaPOD-PMF (base) 0.0562 0.0461 0.0658 0.0488 0.0582 0.0443 0.0721 0.0488 0.0731 0.0589 0.0827 0.0610
Table 7. Performance comparison on top-n recommendation between small and base models.
Methods Sports Beauty Toys
HR@1 HR@5 NDCG@5 HR@10 NDCG@10 HR@1 HR@5 NDCG@5 HR@10 NDCG@10 HR@1 HR@5 NDCG@5 HR@10 NDCG@10
PeaPOD-PMF (small) 0.1252 0.2708 0.2006 0.3596 0.2287 0.1292 0.2628 0.1978 0.3423 0.2228 0.0728 0.2126 0.1581 0.2872 0.1817
PeaPOD-PMF (base) 0.1280 0.2698 0.2016 0.3588 0.2300 0.1000 0.2268 0.1652 0.3086 0.1914 0.0649 0.1494 0.1079 0.2168 0.1291
Table 8. Performance comparison on explanation generation between small and base models.
Methods Sports Beauty Toys
BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
PeaPOD-PMF (small) 0.9148 11.2685 1.7870 8.8078 1.0050 13.4532 1.5863 10.3238 2.4295 13.5537 3.9567 11.3863
PeaPOD-PMF (base) 1.0548 14.1369 2.0106 11.1413 0.8602 13.2299 1.3422 10.0028 2.5037 14.4358 3.9266 11.8846
6.5. Effects of Multi-head Attention
Figure 5. The effects of different number of heads used for attention on the Beauty dataset. Note that ‘0’ represents multi-layer perceptron (MLP) rather than the number of heads.
To understand the effects of multi-head attention versus alternative implementations, we perform an ablation experiment with a varying number of heads. Specifically, to create ablated versions of the collaborative user prompt $\boldsymbol{p}$, we train PeaPOD-PMF on the Beauty dataset with the number of heads set to {1, 2, 4, 8, 16}. In addition, we also experiment with passing the user embedding through a multi-layer perceptron (MLP). Rather than performing attention and passing the attended values through a linear layer, MLP passes the user embedding directly through a linear layer to project the user embedding as a personalized soft prompt. We include this experiment to observe if encoding collaborative information via attention provides superior performance versus simpler methods. We show the experimental results in Figure 5.
We observe that multi-head attention with 4 heads, denoted as $MHA_{4}$, achieved the best overall performance, particularly in sequential recommendation. $MHA_{4}$ showed a slight improvement over single-head attention (referred to as ‘1’ in Figure 5) in all explanation generation metrics. For top-n recommendation, the performance of $MHA_{4}$ and single-head attention was close to equal. We also observe that all attention-based models significantly outperformed the MLP implementation, showing the advantage of the compositional attention strategy. However, increasing the number of heads beyond 4 led to a decline in performance for all three recommendation tasks. This suggests that while adding more heads improves performance up to a certain point, exceeding this threshold causes the model to overfit.
7. Conclusion
In this paper, we propose PeaPOD, a novel generative recommender model that effectively distills individual and group user preferences into a collaborative user prompt. In particular, we introduce a multi-head attention strategy that uses the query-key-value mechanism to generate a personalized user embedding. Multi-head attention allows the model to capture muliple user-to-user relationships and enables the model to efficiently share collaborative knowledge between similar users. Our experiments confirm the effectiveness of our PeaPOD framework in three common recommendation tasks. Future directions include adding additional information during training, such item metadata, and developing an end-to-end architecture to unify the user embedding initialization and multi-head attention stages.
8. Limitations and Ethical Concerns
One limitation of our approach is that we restrict our study to exclusively using user and item IDs. Thus, we do not incorporate any additional metadata such as user profiles or item descriptions. We find that this may hinder performance on explanation generation, which relies on having a strong understanding of both the user and item to provide a personalized response. We predict that providing additional metadata to the model can improve performance, particularly if it is given in natural language text as the LLM is already pre-trained on such data. We plan to extend PeaPOD to utilize this information in future work.
Another concern is the issue of fairness in generative recommender systems. For example, prior research has shown that LLM-based recommenders skew towards popular items (Yoon et al., 2024), which can cause a disparity in recommendations and lead to a bias-amplification loop. In addition, past research has investigated the issue of fairness in the LLM itself (Salutari et al., 2023). Thus, careful consideration must be made when deploying a generative recommender model to mitigate unfair recommendations.